OpenStudy (anonymous):

The graph of y = -3sin( 2x) is shifted to the left by 30 units and up 5 units. Write the new equation.

OpenStudy (anonymous):

If the graph has shifted up 5 units, the graph becomes y = -3sin(2x) + 5 Compare y = sin 2x to y = -3sin (2x) + 5 (forget the 5 for now) shift y = -3 sin (2x) by 30 units to the left, the x-coordinate is -30. The graph now starts x = -30, y = 0 -3sin ([2*-30] = 0 -3sin(-60) = 0 But this is incorrect, therefore we need to add a constant to make it that -3sin(-60 + C) = 0 therefore -3sin(-60+60) = 0 Which is represented as -3sin(2x+60) the final equation is y = 5 - 3 sin (2x + 60)